OpenStudy (anonymous):

Prove the identity: csc^2 x - 1 / cos x = cot x csc x HELP ME PLEAASEEE. :D

terenzreignz (terenzreignz):

What have you got to begin with? Want to start with the Pythagorean?

OpenStudy (anonymous):

okay po. ahm, how po?

terenzreignz (terenzreignz):

With this, to be precise... cos² x + sin² x = 1 Now, if you divide both sides of the equation by sin² x, you get cot² x + 1 = csc² x

terenzreignz (terenzreignz):

Please give me some indication that you are able to follow.

OpenStudy (anonymous):

cot² x + 1 = csc² x uuse this formula

OpenStudy (anonymous):

what po? i didn't get it po. sorry, im so slow po.

terenzreignz (terenzreignz):

You know the Pythagorean identity cos² x + sin² x = 1 ?

OpenStudy (anonymous):

@terenzreignz u can't use rhs side since here u hav to prove .did u gt wat i m saying

OpenStudy (anonymous):

yes po .. i know that po ..

terenzreignz (terenzreignz):

@manishsatywali I'm not using the rhs, I'm deriving the formula you just gave.

OpenStudy (anonymous):

k pardon me............

terenzreignz (terenzreignz):

Okay, well, starting from cos² x + sin² x = 1 Well, divide both sides by sin² x, you get \[\frac{\cos^{2}x}{\sin^{2}x}+\frac{\sin^{2}x}{\sin^{2}x}=\frac{1}{\sin^{2}x}\] Right?

OpenStudy (anonymous):

yes po. then po?

terenzreignz (terenzreignz):

Okay, we can bring out the exponents, right? \[\left( \frac{\cos x}{\sin x} \right)^{2}+1=\left( \frac{1}{\sin x}\right)^{2}\] Okay?

OpenStudy (anonymous):

okay po . then after that po?

terenzreignz (terenzreignz):

What's cos x / sin x ?

OpenStudy (anonymous):

cot x po .

terenzreignz (terenzreignz):

Very good. Now what's 1/sin x ?

OpenStudy (anonymous):

csc x po .

terenzreignz (terenzreignz):

So we can rewrite that equation as cot² x + 1 = csc² x am I right?

OpenStudy (anonymous):

is it on the right side?

terenzreignz (terenzreignz):

What's on the right side? I just replaced cos x / sin x with cot x and 1/sin x with csc x

OpenStudy (anonymous):

ahh .. okays po . i get it . then po?

terenzreignz (terenzreignz):

Oh, we're working on the left side, for now...

OpenStudy (anonymous):

okay po ..

terenzreignz (terenzreignz):

Okay, well, let me rewrite it cot² x + 1 = csc² x Then we can subtract 1 from both sides, and get cot² x = csc² x - 1 Are you catching me so far?

OpenStudy (anonymous):

yes i am ..

terenzreignz (terenzreignz):

Well, then, you have (csc² x - 1)/cos x On the left side, right? Slowly, we'll turn it into the right side... First off, you know for a fact that csc² x - 1 is equal to...?

OpenStudy (anonymous):

cot^2 x po .

terenzreignz (terenzreignz):

So we can turn the left side into (cot² x) / cos x right?

OpenStudy (anonymous):

yes po.

terenzreignz (terenzreignz):

Let me just write it more neatly: \[\frac{\cot^2 x}{\cos x}=(\cot^2 x)\left( \frac{1}{\cos x} \right)\] Getting it so far?

OpenStudy (anonymous):

yes po .. you separate the two ..

terenzreignz (terenzreignz):

Well, we're going to separate a little more, in particular, let's express cot x as cos x / sin x shall we? \[=\left(\frac{\cos x}{\sin x}\right)\left(\frac{\cos x}{\sin x}\right)\left(\frac{1}{\cos x}\right)\] Saw how I did that? Remember that it was cot² x, originally...

OpenStudy (anonymous):

aha! i get it. :D you separate cot^2 x in a simplest form ..

terenzreignz (terenzreignz):

Actually, that wasn't necessary... I can replace one of them with cot x again. cos x / sin x = cot x, right? Well \[=(\cot x)\left(\frac{\cos x}{\sin x}\right)\left(\frac{1}{\cos x}\right)\] okay?

OpenStudy (anonymous):

aha! you're right. then po?

terenzreignz (terenzreignz):

LOL Anyway... You can see that you can cancel out the cos x in the expression, right?

OpenStudy (anonymous):

yes po. hehe

terenzreignz (terenzreignz):

Well, cancel out and you get \[(\cot x)\left( \frac{1}{\sin x}\right)\] Can you see it now? What's 1/sin x ?

OpenStudy (anonymous):

csc x po. haha. i know the answer now. thank you po. :D

terenzreignz (terenzreignz):

heh